# Lattice determination of the QCD low-energy constant $\ell_{\scriptscriptstyle{7}}$

### v1

@@ III.4 Chiral extrapolation of $\ell_{\scriptscriptstyle{7}}$
0.3197 6.66(51)stat(15)syst 384(7) 129.0(2.3) 0.432(40) 0.5541(42) 1.59(19)stat(04)syst
0.2131 5.17(53)stat(07)syst 316(6) 122.6(2.3) 0.494(47) 0.6125(37) 1.56(22)stat(03)syst
0.1421 4.58(45)stat(29)syst 265(6) 115.1(3.1) 0.625(71) 0.6715(34) 1.92(29)stat(12)syst
Table 3: Continuum limits of $\ell_{\scriptscriptstyle{7}}$ as a function of the light-to-strange quark mass ratio $R$. We also report the continuum limits of $M_{\pi}(R)$ and $F_{\pi}(R)$ obtained in [15] (their reported errors include both statistical and systematic uncertainties), and of the improved estimator $\tilde{\ell}_{\scriptscriptstyle{7}}(R)$ where leading logs have been removed via $c_{1}(R)$ and $c_{2}(R)$ (see Sec. III.3 for more details).
@@
The calculation of the strong pion mass splitting at NNLO in $\chi\mathrm{PT}$ is fully carried out for the first time in this study. In the following, we will report the results that are necessary for the chiral extrapolation of $\ell_{\scriptscriptstyle{7}}$, while further details can be found in Appendix E.
Remarkably, all NNLO contributions to the pion mass splitting are $\mathcal{O}\left[(\Delta m)^{2}\right]$, and their structure turns out to be rather simple, as there are only two $\mathcal{O}(m_{\scriptscriptstyle{\ell}})$ terms appearing in Eq. (35). One is a chiral logarithm of order $\mathcal{O}(m_{\scriptscriptstyle{\ell}}\log m_{\scriptscriptstyle{\ell}})$, which turns out to be proportional to $\ell_{\scriptscriptstyle{7}}$ itself—see also [35, 36], where the calculation of the leading non-analytic contributions to $\Delta M_{\pi}^{2}$, once matched to our notation, yields a chiral log multiplying $\ell_{\scriptscriptstyle{7}}$ with the same coefficient as ours. The second is a linear $\mathcal{O}(m_{\scriptscriptstyle{\ell}})$ correction, whose coefficient depends instead on some combination of the couplings of the NNLO isospin-breaking operators appearing at $\mathcal{O}(p^{6})$, see Appendix E. All other corrections are higher-order in the chiral expansion, i.e., $\mathcal{O}(m_{\scriptscriptstyle{\ell}}^{2})$ in the quark mass. In the end, Eqs. (1) and (10) are modified respectively as follows:
@@
The first factor, $c_{1}(R)$, takes into account the fact that, for the purpose of evaluating the pre-factor of $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}$ in Eq. (23) we have used the simulated values of $M_{\pi}^{2}(R)/m_{\scriptscriptstyle{\ell}}$ and $F_{\pi}(R)$, cf. Tab. 1, rather than their expression in the chiral limit $M^{2}/m_{\scriptscriptstyle{\ell}}=2B$ and $F$, which is what actually appears in the left-hand side of Eq. (1), see [36, 2, 3]. This choice is irrelevant for the purpose of computing the chiral limit of $\ell_{\scriptscriptstyle{7}}(R)$, but of course introduces additional finite-quark-mass corrections to Eq. (36), that $c_{1}(R)$ aims to cancel. The second factor $c_{2}(R)$ aims instead to remove the $\mathcal{O}(m_{\scriptscriptstyle{\ell}}\log m_{\scriptscriptstyle{\ell}})$ leading-log term appearing in the right-hand side of Eq. (36). To compute $c_{1}$ and $c_{2}$, we took $F$ from the latest FLAG average, $F=86.6(6)$ MeV [5], and $M^{2}/R=Bm_{\scriptscriptstyle{\rm s}}=[2.36(5)\times 10^{5}]~\mathrm{MeV}^{2}$ from [15]. For the finite-quark-mass quantities $M_{\pi}(R)$ and $F_{\pi}(R)$, instead, we took the continuum-extrapolated results of [15], reported in Tab. 3. The correction factors and the corrected quantity $\tilde{\ell}_{\scriptscriptstyle{7}}$ are found in Tab. 3.
@@
i.e., according to a leading linear dependence in $R$, with no $\mathcal{O}(R\log R)$ terms, up to $\mathcal{O}(R^{2})$ sub-leading terms. As it can be seen from Fig. 4, the corrected data show almost a flat behavior as a function of $R$, $\tilde{C}\simeq(-1.2\pm 1.8)\cdot 10^{-3}$, and are thus in perfect agreement with our expectations. A naive linear extrapolation in $R$ of the uncorrected data would still give a marginally compatible extrapolation, but with a much larger slope.
In order to propagate systematic uncertainties associated with the continuum limit to the chiral extrapolation, we followed a bootstrap-based strategy. In particular, we performed a bootstrap resampling of each data point, and added to each value of $\ell_{\scriptscriptstyle{7}}(R)$ a zero-mean stochastic Gaussian noise with variance equal to the squared systematic error. For each bootstrap sample we repeated the chiral fit, and estimated the final systematic error from the average absolute deviation between the central value of the chiral limits of each bootstrap sample and the central value of the chiral extrapolation obtained ignoring systematic errors.
Concerning instead the systematic error related to the possible influence of higher-order $\mathcal{O}(R^{2})$ terms, with only three points it is not possible to perform a comprehensive study. A possible crude estimate can however be obtained, using again Eq. (33), by comparing the linear extrapolation presented above with the one obtained from a 0-degrees-of-freedom best fit to the data corresponding to the two smallest values of $R$. This uncertainty is added in quadrature to the one obtained from the propagation of the systematic errors on the continuum limits.
@@
$\ell_{\scriptscriptstyle{7}}\times 10^{3}=1.98(48)_{\mathrm{stat}}(26)_{\mathrm{syst}}=1.98(54)_{\mathrm{tot}}.$ (41)
As it can be seen, our final error is dominated by statistical uncertainties (mainly due to the chiral extrapolation), which account for about the $85\%$ of the total error budget. Therefore, further reducing it would require determinations of $\ell_{\scriptscriptstyle{7}}$ for lower quark masses. Unfortunately, the signal-to-noise ratio of $\ell_{\scriptscriptstyle{7}}$ rapidly deteriorates as we approach the physical point, mainly due to the disconnected contribution to $\ell_{\scriptscriptstyle{7}}$ (see Appendix D for more details). Thus, determining $\ell_{\scriptscriptstyle{7}}$ at, for example, the physical point $R\simeq 0.0355$ would require a very challenging numerical effort.
@@ IV Conclusions
$\ell_{\scriptscriptstyle{7}}\times 10^{3}=1.98(48)_{\mathrm{stat}}(26)_{\mathrm{syst}}=1.98(54)_{\mathrm{tot}}.$ (42)
Our result is compatible with, and significantly improves on, previous lattice determinations, $\ell_{\scriptscriptstyle{7}}\times 10^{3}=6.5(3.8)$ [6] (indirect determination), $\ell_{\scriptscriptstyle{7}}\times 10^{3}=2.5(1.4)$ [11] (direct determination for a single gauge ensemble), as well as with the phenomenological estimate $\ell_{\scriptscriptstyle{7}}\times 10^{3}=7(4)$ [4].
@@ Acknowledgements.
It is a pleasure to thank C. Bonati for useful discussions, and M. Gorghetto and G. Villadoro for communications about Ref. [7]. The work of C. Bonanno is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grant PID2021-127526NB-I00, both of which are funded by MCIN/AEI/10.13039/501100011033. R. Dionisio acknowledges funding from the European Union Next-Generation EU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 1, INVESTMENT N.4.1 – CUP N. I51J24000160007 (PhD Cycle XL, Ministerial Decree no. 629/2024). This work has been supported by the project “Non-perturbative aspects of fundamental interactions, in the Standard Model and beyond” funded by MUR, Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN), Bando 2022, grant 2022TJFCYB (CUP I53D23001440006). Numerical calculations have been performed on the Leonardo machine at Cineca, based on the agreement between INFN and Cineca, under projects INF23_npqcd and INF24_npqcd.
@@ Appendix A Taste symmetry restoration in the continuum limit of the pion mass
Appendix B Comparison of the disconnected diagrams in the pseudo-scalar and axial channels
@@
Appendix C Determination of $\ell_{\scriptscriptstyle{7}}$ from the vector channel and comparison with the axial-vector one
As stated in Sec. III, the $\ell_{\scriptscriptstyle{7}}$ estimator in Eq. (23) was computed using the axial-vector pion interpolating operator, as no signal was observed using the customary pseudoscalar channel. As a further check of our procedure, we also extracted $\ell_{\scriptscriptstyle{7}}$ using Eq. (23) from the vector channel. Despite results differing from the axial-vector ones at finite lattice spacing, the two determinations nicely converge towards the same continuum limit, see Fig. 7. This is not surprising, giving that all interpolating operators, in the continuum limit, describe exactly the same particle with the same mass, cf. also Appendix A. Since the axial-vector channel gave smaller uncertainties with respect to the vector one, we chose the former to perform our final calculations.
@@
Appendix D Scaling of the signal-to-noise ratio of $\ell_{\scriptscriptstyle{7}}$ towards the chiral limit
@@
Appendix E Charged and neutral pion masses at NNLO in $\chi\mathrm{PT}$
@@
The expressions for $m_{1}$ and $\delta_{1}$ date back to Gasser and Leutwyler [36, 2, 3], while the expression for $m_{2}$ was previously computed in Refs. [38, 39] only in the isospin-symmetric case $\Delta=0$, and neglecting the contributions coming from the NNLO LECs, as these were only introduced later [40, 34]. Thus, the full expression of $m_{2}$ (including isospin-breaking corrections and the NNLO LECs contributions) and of $\delta_{2}$ are new results first presented in this study. Note also that our NNLO formula for $M_{\pi^{0}}^{2}$ corrects a misprint in Eq. (A6) of Ref. [7] [^2] As confirmed by the authors of [7] via private communication..
@@
$\frac{\Lambda_{{\scriptscriptstyle{\rm IB}}}}{\mu_{0}}=\frac{\Lambda_{{\scriptscriptstyle{\rm IB}}}}{4\pi F}=0.83(33).$ (61)

### v2

@@ III.4 Chiral extrapolation of $\ell_{\scriptscriptstyle{7}}$
0.3197 6.66(51)stat(15)syst 384(7) 105.1(2.9) 0.637(61) 0.5541(42) 2.37(29)stat(04)syst
0.2131 5.17(53)stat(07)syst 316(6) 100.7(2.8) 0.762(52) 0.6125(37) 2.42(27)stat(03)syst
0.1421 4.58(45)stat(29)syst 265(6) 97.7(3.1) 0.857(61) 0.6715(34) 2.64(32)stat(12)syst
Table 3: Continuum limits of $\ell_{\scriptscriptstyle{7}}$ as a function of the light-to-strange quark mass ratio $R$. We also report the continuum limits of $M_{\pi}(R)$ and $F_{\pi}(R)$ (reported errors include both statistical and systematic uncertainties), and of the improved estimator $\tilde{\ell}_{\scriptscriptstyle{7}}(R)$ where leading logs have been removed via $c_{1}(R)$ and $c_{2}(R)$ (see Sec. III.3 for more details).
@@
The calculation of the strong pion mass splitting at NNLO in $\chi\mathrm{PT}$ is fully carried out for the first time in this study. In the following, we will report the results that are necessary for the chiral extrapolation of $\ell_{\scriptscriptstyle{7}}$, while further details can be found in Appendix D.
Remarkably, all NNLO contributions to the pion mass splitting are $\mathcal{O}\left[(\Delta m)^{2}\right]$, and their structure turns out to be rather simple, as there are only two $\mathcal{O}(m_{\scriptscriptstyle{\ell}})$ terms appearing in Eq. (35). One is a chiral logarithm of order $\mathcal{O}(m_{\scriptscriptstyle{\ell}}\log m_{\scriptscriptstyle{\ell}})$, which turns out to be proportional to $\ell_{\scriptscriptstyle{7}}$ itself—see also [35, 36], where the calculation of the leading non-analytic contributions to $\Delta M_{\pi}^{2}$, once matched to our notation, yields a chiral log multiplying $\ell_{\scriptscriptstyle{7}}$ with the same coefficient as ours. The second is a linear $\mathcal{O}(m_{\scriptscriptstyle{\ell}})$ correction, whose coefficient depends instead on some combination of the couplings of the NNLO isospin-breaking operators appearing at $\mathcal{O}(p^{6})$, see Appendix D. All other corrections are higher-order in the chiral expansion, i.e., $\mathcal{O}(m_{\scriptscriptstyle{\ell}}^{2})$ in the quark mass. In the end, Eqs. (1) and (10) are modified respectively as follows:
@@
The first factor, $c_{1}(R)$, takes into account the fact that, for the purpose of evaluating the pre-factor of $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}$ in Eq. (23) we have used the simulated values of $M_{\pi}^{2}(R)/m_{\scriptscriptstyle{\ell}}$ and $F_{\pi}(R)$, cf. Tab. 1, rather than their expression in the chiral limit $M^{2}/m_{\scriptscriptstyle{\ell}}=2B$ and $F$, which is what actually appears in the left-hand side of Eq. (1), see [36, 2, 3]. This choice is irrelevant for the purpose of computing the chiral limit of $\ell_{\scriptscriptstyle{7}}(R)$, but of course introduces additional finite-quark-mass corrections to Eq. (36), that $c_{1}(R)$ aims to cancel. The second factor $c_{2}(R)$ aims instead to remove the $\mathcal{O}(m_{\scriptscriptstyle{\ell}}\log m_{\scriptscriptstyle{\ell}})$ leading-log term appearing in the right-hand side of Eq. (36). To compute $c_{1}$ and $c_{2}$, we took $F$ from the latest FLAG average, $F=86.6(6)$ MeV [5], and $M^{2}/R=Bm_{\scriptscriptstyle{\rm s}}=[2.36(5)\times 10^{5}]~\mathrm{MeV}^{2}$ from [15]. For the finite-quark-mass quantities $M_{\pi}(R)$ and $F_{\pi}(R)$, instead, we took the continuum-extrapolated results of [15], reported in Tab. 3. The correction factors and the corrected quantity $\tilde{\ell}_{\scriptscriptstyle{7}}$ are also found in Tab. 3.
@@
i.e., according to a leading linear dependence in $R$, with no $\mathcal{O}(R\log R)$ terms, up to $\mathcal{O}(R^{2})$ sub-leading terms. As it can be seen from Fig. 4, the corrected data show almost a flat behavior as a function of $R$, $\tilde{C}\simeq(-1.4\pm 2.4)\cdot 10^{-3}$, and are thus in perfect agreement with our expectations. A naive linear extrapolation in $R$ of the uncorrected data would still give a compatible extrapolation, but with a much larger slope.
Figure 5: Comparison of our $N_{\rm f}=2+1$ lattice determination of $\ell_{\scriptscriptstyle{7}}$ (starred point, shaded area) with previous ones from Refs. [6, 11], and with the phenomenological estimate of [4].
In order to propagate systematic uncertainties associated with the continuum limit to the chiral extrapolation, we followed a bootstrap-based strategy. In particular, we performed a bootstrap resampling of each data point, and added to each value of $\ell_{\scriptscriptstyle{7}}(R)$ a zero-mean stochastic Gaussian noise with variance equal to the squared systematic error. For each bootstrap sample we repeated the chiral fit, and estimated the final systematic error from the average absolute deviation between the central value of the chiral limits of each bootstrap sample and the central value of the chiral extrapolation obtained ignoring systematic errors. Concerning instead the systematic error related to the possible influence of higher-order $\mathcal{O}(R^{2})$ terms, with only three points it is not possible to perform a comprehensive study. A possible crude estimate can however be obtained, using again Eq. (33), by comparing the linear extrapolation presented above with the one obtained from a 0-degrees-of-freedom best fit to the data corresponding to the two smallest values of $R$. This uncertainty is added in quadrature to the one obtained from the propagation of the systematic errors on the continuum limits.
@@
$\ell_{\scriptscriptstyle{7}}\times 10^{3}=2.79(58)_{\mathrm{stat}}(19)_{\mathrm{syst}}=2.79(61)_{\mathrm{tot}}.$ (41)
As it can be seen, our final error is dominated by statistical uncertainties (mainly due to the chiral extrapolation), which account for about the $85\%$ of the total error budget. Therefore, further reducing it would require determinations of $\ell_{\scriptscriptstyle{7}}$ for lower quark masses. Unfortunately, the signal-to-noise ratio of $\ell_{\scriptscriptstyle{7}}$ rapidly deteriorates as we approach the physical point, mainly due to the disconnected contribution to $\ell_{\scriptscriptstyle{7}}$ (see Appendix C for more details). Thus, determining $\ell_{\scriptscriptstyle{7}}$ at, for example, the physical point $R\simeq 0.0355$ would require a very challenging numerical effort.
@@ IV Conclusions
$\ell_{\scriptscriptstyle{7}}\times 10^{3}=2.79(58)_{\mathrm{stat}}(19)_{\mathrm{syst}}=2.79(61)_{\mathrm{tot}}.$ (42)
A comparison with previous estimates is displayed in Fig. 5. Our result is compatible with, and significantly improves on, previous lattice determinations, $\ell_{\scriptscriptstyle{7}}\times 10^{3}=6.5(3.8)_{\rm stat}(0.20)_{\rm syst}$ [6] (indirect determination), $\ell_{\scriptscriptstyle{7}}\times 10^{3}=2.50(1.30)_{\rm stat}(0.50)_{\rm syst}$ [11] (direct determination for a single gauge ensemble), as well as with the phenomenological estimate $\ell_{\scriptscriptstyle{7}}\times 10^{3}=7(4)$ [4].
@@ Acknowledgements.
It is a pleasure to thank C. Bonati for useful discussions, and M. Gorghetto and G. Villadoro for communications about Ref. [7]. The work of C. Bonanno is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grant PID2021-127526NB-I00, both of which are funded by MCIN/AEI/10.13039/501100011033. R. Dionisio acknowledges funding from the European Union (EU) Next Generation EU (NGEU) – National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 1, INVESTMENT N.4.1 – CUP N. I51J24000160007 (PhD Cycle XL, Ministerial Decree no. 629/2024). F. Sanfilippo is supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by the EU NGEU – and by the Italian Ministry of University and Research (MUR) project FIS 00001556. This work has also been supported by the project “Non-perturbative aspects of fundamental interactions, in the Standard Model and beyond” funded by MUR, Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN), Bando 2022, grant 2022TJFCYB (CUP I53D23001440006). Numerical calculations have been performed on the Leonardo machine at Cineca, based on the agreement between INFN and Cineca, under projects INF23_npqcd and INF24_npqcd.
@@ Appendix A Taste symmetry restoration in the continuum limit of the pion mass
Appendix B Comparison of $\ell_{\scriptscriptstyle{7}}$ determinations from the pseudo-scalar, vector and axial-vector channels
@@
As a further check of our procedure, we also extracted $\ell_{\scriptscriptstyle{7}}$ using Eq. (23) from the vector channel. Despite results differing from the axial-vector ones at finite lattice spacing, the two determinations nicely converge towards the same continuum limit, see Fig. 8. This is not surprising, giving that all interpolating operators, in the continuum limit, describe exactly the same particle with the same mass, cf. also Appendix A. Since the axial-vector channel gave smaller uncertainties with respect to the vector one, we chose the former to perform our final calculations.
@@
Appendix C Scaling of the signal-to-noise ratio of $\ell_{\scriptscriptstyle{7}}$ towards the chiral limit
@@
Appendix D Charged and neutral pion masses at NNLO in $\chi\mathrm{PT}$
@@
The expressions for $m_{1}$ and $\delta_{1}$ date back to Gasser and Leutwyler [36, 2, 3], while the expression for $m_{2}$ was previously computed in Refs. [38, 39] only in the isospin-symmetric case $\Delta=0$, and neglecting the contributions coming from the NNLO LECs, as these were only introduced later [40, 34]. Thus, the full expression of $m_{2}$ (including isospin-breaking corrections and the NNLO LECs contributions) and of $\delta_{2}$ are new results first presented in this study. Note also that our NNLO formula for $M_{\pi^{0}}^{2}$ corrects a misprint in Eq. (A6) of Ref. [7]. [^4] As confirmed by the authors of [7] via private communication.
@@
$\frac{\Lambda_{{\scriptscriptstyle{\rm IB}}}}{\mu_{0}}$ $=$ $\frac{\Lambda_{{\scriptscriptstyle{\rm IB}}}}{4\pi F}=0.87(32),$ (61)
$\implies\Lambda_{{\scriptscriptstyle{\rm IB}}}$ $\simeq$ $0.95(35)\text{ GeV}.$ (62)
Appendix E Determination of $\bar{\ell}_{\scriptscriptstyle{3}}$ and $\bar{\ell}_{\scriptscriptstyle{4}}$
[…]
Figure 11: Comparison of our $N_{\rm f}=2+1$ lattice determinations of $\bar{\ell}_{\scriptscriptstyle{3}}$ (top panel) and $\bar{\ell}_{\scriptscriptstyle{4}}$ (bottom panel), drawn as starred points, with previous ones [41, 42, 43, 44, 6] entering FLAG 21 averages [5] (square points and shaded bands). We also show the phenomenological estimates of [2, 45].
The determinations of $M_{\pi}$ and $F_{\pi}$ reported in the present study in Tab. 3, and used for the calculation of $\ell_{\scriptscriptstyle{7}}$, can also be employed to extract two other NLO LECs: $\bar{\ell}_{\scriptscriptstyle{3}}$ and $\bar{\ell}_{\scriptscriptstyle{4}}$. These couplings parameterize the corrections to the chiral limits of the pion mass and of the pion decay constant [3]:
$M_{\pi}^{2}$ $=$ $M^{2}\left\{1-\frac{M^{2}}{32\pi^{2}F^{2}}\left[\bar{\ell}_{\scriptscriptstyle{3}}-L(M)\right]\right\},$ (63)
$F_{\pi}$ $=$ $F\left\{1+\frac{M^{2}}{16\pi^{2}F^{2}}\left[\bar{\ell}_{\scriptscriptstyle{4}}-L(M)\right]\right\},$ (64)
with
$L(M)$ $=$ $\log\left(\frac{M^{2}}{{M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}}^{2}}\right),$ (65)
$M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}$ $=$ $135\text{ MeV},$ (66)
a chiral log. Using $M^{2}=2Bm_{\scriptscriptstyle{\ell}}=2Bm_{\scriptscriptstyle{\rm s}}R$, one can perform a four-parameter global fit of $M_{\pi}(R)$ and $F_{\pi}(R)$ and extract $\bar{\ell}_{\scriptscriptstyle{3}}$ and $\bar{\ell}_{\scriptscriptstyle{4}}$:
$\frac{M_{\pi}^{2}(R)}{R}$ $=$ $2Bm_{\scriptscriptstyle{\rm s}}\left\{1-\frac{Bm_{\scriptscriptstyle{\rm s}}R}{16\pi^{2}F^{2}}\left[\bar{\ell}_{\scriptscriptstyle{3}}-\tilde{L}(R)\right]\right\},$ (67)
$F_{\pi}(R)$ $=$ $F\left\{1+\frac{Bm_{\scriptscriptstyle{\rm s}}R}{8\pi^{2}F^{2}}\left[\bar{\ell}_{\scriptscriptstyle{4}}-\tilde{L}(R)\right]\right\},$ (68)
where now the chiral log reads
$\tilde{L}(R)=\log\left(\frac{2Bm_{\scriptscriptstyle{\rm s}}R}{{M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}}^{2}}\right).$ (69)
The joint fit was performed after the application of the finite-volume corrections $R_{M_{\pi}}$ and $R_{F_{\pi}}$ (cf. Sec. III.2), which turned out to be of always smaller than our statistical uncertainties.
The LO LECs $\Sigma m_{\scriptscriptstyle{\rm s}}=Bm_{\scriptscriptstyle{\rm s}}\times F^{2}$ and $F$ were already given in the original paper [15], while $\bar{\ell}_{\scriptscriptstyle{3}}$ and $\bar{\ell}_{\scriptscriptstyle{4}}$ are reported here for the first time. We find:
$\bar{\ell}_{\scriptscriptstyle{3}}$ $=$ $4.18(60)_{\rm stat}(70)_{\rm syst}=4.18(92)_{\rm tot},$ (70)
$\bar{\ell}_{\scriptscriptstyle{4}}$ $=$ $3.90(20)_{\rm stat}(27)_{\rm syst}=3.90(34)_{\rm tot}.$ (71)
The analysis for $\bar{\ell}_{\scriptscriptstyle{3}}$ and $\bar{\ell}_{\scriptscriptstyle{4}}$ goes along the same lines of the one for $\ell_{\scriptscriptstyle{7}}$ presented in the main text. The central values and the statistical uncertainties are obtained from the joint chiral fit including all available points, shown in Fig. 10. The systematic error is the result of the quadrature sum of two error sources. One is the propagation of the systematic uncertainties on $F_{\pi}$ and $M_{\pi}$ due to the continuum extrapolation. The other is the observed difference in the fit results for $\bar{\ell}_{\scriptscriptstyle{3}}$ and $\bar{\ell}_{\scriptscriptstyle{4}}$ when including/excluding the point at the heaviest pion mass from the joint chiral fit.
Our chiral extrapolations of $F_{\pi}$ and $M_{\pi}$ are well under control, as it can be seen by the very good agreement of our results in $R=R^{{\scriptscriptstyle{(\mathrm{phys})}}}\simeq 0.036$ [5] with the physical values $M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}=135$ MeV and $F_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}=92$ MeV. Moreover, our results for $\bar{\ell}_{\scriptscriptstyle{3}}$ and $\bar{\ell}_{\scriptscriptstyle{4}}$ are in very good agreement both with the latest FLAG world averages [5] and with previous determinations in the literature. This comparison is illustrated in Fig. 11.